The beauty of comparative advantage is that if you have a lower opportunity cost in the production of one good or service versus another, and your potential trading partner has the opposite position, you’re both better off in the transaction. True or False? Explain.
Answer #1

TRUE

Let us understand this with a small and simple example :

Let country A and B manufacture two goods : Rice and Car .

Country A can make 5 Rice or 10 Cars with its resources .

Country B can make 10 Rice or 15 Cars .

Now , country A's opportunity cost of making 1 Rice = 10/5 = 2 Cars

Similarly for Country B = 15/10 = 1.5 Cars .

Hence , we can see that B has lower opportunity cost of making Rice , hence has a comparative advantage in making Rice . So country B will specialize in rice . Vice versa country A has comparative advantage in Cars ( opportunity cost of A = 0.5 Rice , B = 0.67 Rice ) . So A will specialize in Cars .

New production :

A = 10 Cars . B = 10 Rice .

If they trade at 1 Rice = 1 Car .

A can give 5 cars and get 5 rice .

B can get 5 cars and give 5 rice .

Before trade if A made 5 cars then car consumption would be 0 . So there is gains from trade .